Traveling at a rate of speed of 70 miles per hour. Traveling 125 miles what is my expected arrival time equal
Answer by rwm(914)

r*t=d
d/r=t
125/70=t
1 55/70 hrs
If you don't get stopped by the state police or have an accident caused by excessive speed.